### v3

# MIBoost: A Gradient Boosting Algorithm for Variable Selection After Multiple Imputation

Abstract
Statistical learning methods for automated variable selection, such as LASSO, elastic nets, or gradient boosting, have become increasingly popular tools for building powerful prediction models. Yet, in practice, analyses are often complicated by missing data. The most widely used approach to address missingness is multiple imputation, which involves creating several completed datasets. However, there is an ongoing debate on how to perform model selection in the presence of multiple imputed datasets. Simple strategies, such as pooling models across datasets, have been shown to have suboptimal properties. Although more sophisticated methods exist, they are often difficult to implement and therefore not widely applied. In contrast, two recent approaches modify the regularization methods LASSO and elastic nets by defining a single loss function, resulting in a unified set of coefficients across imputations. Our key contribution is to extend this principle to the framework of component-wise gradient boosting by proposing MIBoost, a novel algorithm that employs a uniform variable-selection mechanism across imputed datasets. Our approach is implemented in the form of the R package booami.
Keywords: Gradient boosting, Multiple imputation, Variable selection, Component‑wise gradient boosting, Prediction
1 Introduction
Multiple imputation is widely considered the gold standard for handling missing data. When the main interest lies in causal inference, coefficients and standard errors obtained from each imputed dataset can be easily aggregated using Rubin’s pooling rules [1]. However, performing variable selection for prediction purposes with multiple imputed datasets remains methodologically challenging, as the selected models can differ substantially across datasets. Many previously proposed methods for obtaining a single unified model are complex and insufficiently implemented in software packages. Simply selecting all variables chosen in any of the datasets often leads to overly complex models [2]; hence, many users resort to ad hoc solutions, such as including only variables selected in a minimum proportion of (e.g., 50%) of imputed datasets. While this heuristic may yield satisfactory results in some cases, it suffers from a strong dependence on an arbitrary threshold and can be unstable in small samples or in the presence of correlated predictors [3]. To address this problem, [4] proposed a framework that leverages the popular variable selection methods LASSO and elastic nets [5]. They propose a framework consisting of two different approaches, both of which minimize a single loss function, ultimately yielding a unified model. This framework is implemented in the R package miselect [6]. Another statistical-learning method often considered an alternative to LASSO and elastic nets is component-wise gradient boosting [7], which implicitly results in coefficient shrinkage and variable selection, too, making it well-suited for high-dimensional data. Whereas LASSO and elastic nets are limited to generalized linear models, boosting accommodates diverse loss functions and nonlinear effects, making it highly versatile. However, no method to date ensures uniform variable selection when applying component-wise gradient boosting to multiply imputed data. This paper extends the idea of defining a single loss function across imputations and proposes a novel algorithm for component-wise gradient boosting that enables consistent variable selection across all imputed datasets. This approach is implemented in the R package booami to enable accessibility for methodologically oriented as well as application-oriented researchers.
Our main contributions are as follows:
  • We propose MIBoost, a novel boosting algorithm specifically designed for prediction after multiple imputation, which couples base-learner selection across imputed datasets to improve model stability and interpretability.
  • We develop a cross-validation strategy for MIBoost that avoids data leakage by separating training and validation imputations and aggregates performance across imputations in a principled way.
  • We conduct a comprehensive simulation study assessing the predictive accuracy, variable-selection performance, and computational efficiency of MIBoost compared to existing boosting-based and imputation-based approaches.
  • We provide an R package, booami [8], that implements our proposed approach.
The remainder of this paper is structured as follows: Section 2 provides background on existing methods for variable selection after multiple imputation and on component‑wise gradient boosting. Section 3 introduces the proposed MIBoost algorithm together with its cross‑validation framework, MIBoostCV. Section 4 presents a comprehensive simulation study comparing the performance of MIBoost with established methods. Finally, Section 5 discusses our findings and outlines future research directions.
2 Background
2.1 Statistical and Component-Wise Gradient Boosting
Statistical gradient boosting [9] is a machine learning algorithm that minimizes a specified differentiable loss function $\rho$ by incrementally constructing an additive model in a stagewise manner [10, 11]. Given $n$ observations, we define the empirical loss function as
$\hat{R}(\bm{\eta})=\sum_{i=1}^{n}\rho\big(y_{i},\eta(\bm{X}_{i})\big),$
[…]
$\bm{u}^{[t]}=-\nabla_{\bm{\eta}}\hat{R}(\bm{\eta}),\qquad\bm{u}^{[t]}\in\mathbb{R}^{n}.$
which corresponds to the negative gradient of the empirical loss with respect to the additive predictor, evaluated at its current value (often referred to as pseudo-residuals). In the special case of squared error loss,
$\rho(y,\eta)=\tfrac{1}{2}(y-\eta)^{2},$
these pseudo-residuals coincide with the ordinary residuals.
$\bm{u}=\bm{y}-\hat{\bm{\eta}},\qquad\bm{u}\in\mathbb{R}^{n}.$
Design Matrix and Base-learners
The additive predictor $\hat{\bm{\eta}}$ is a function of the design matrix $\bm{X}$, which is constructed from a set of $c$ raw covariates (e.g., age, sex). From these covariates, we derive a collection of $p$ candidate covariates ($c\leq p$), which form the columns of the design matrix,
$\bm{X}=[\bm{X}_{1},\dots,\bm{X}_{p}]=\begin{bmatrix}\bm{x}_{1}^{T}\\ \vdots\\ \bm{x}_{n}^{T}\end{bmatrix}$,
where $\bm{X}_{j}$ denotes the vector of values of the $j$-th candidate covariate across all observations, and $\bm{x}_{i}$ the $i$-th row vector containing all candidate covariate values for observation $i$. For convenience, we use the shorthand $\eta_{i}:=\eta(\bm{x}_{i})$ for the additive predictor of observation $i$. Candidate covariates may be simple transformations (e.g., age, age2) or more flexible effects, such as penalized splines or random intercepts. To ensure identifiability and comparability, candidate covariates (i.e., their design matrices) are typically centered so that their contribution is orthogonal to the intercept. Building on these candidate covariates, we define base-learners as simple prediction rules (e.g., univariate regressions or smoothers) that estimate the association of each candidate covariate with the outcome.
In this paper we focus on component-wise gradient boosting [7], in which each base-learner corresponds only to a single candidate covariate. Furthermore, at each iteration, only a single base-learner is selected and updated, while all others remain unchanged. Formally, at each boosting iteration, denoted by $t$ in the following, we compute, for each of the $p$ base-learners, the least-squares projection of the current negative gradient vector $\bm{u}^{[t]}$ (the pseudo-residuals) onto the function space of that base-learner:
$\hat{h}_{\beta_{r}}^{[t]}=\arg\min_{h\in\mathcal{H}_{r}}\sum_{i=1}^{n}\big(u_{i}^{[t]}-h(x_{i,r})\big)^{2},\qquad\hat{h}_{\beta_{r}}^{[t]}:\mathbb{R}\to\mathbb{R}.$
where $\mathcal{H}_{r}$ denotes the function space associated with base-learner $r$. The estimated function $\hat{h}_{\beta_{r}}^{[t]}$ is referred to as the empirical base-learner for candidate covariate $r$ in iteration $t$.
For a simple linear base-learner, this corresponds to an ordinary least-squares regression of the pseudo-residuals on an intercept and the $r$-th covariate. This yields the fitted function
$\hat{h}_{\beta_{r}}^{[t]}(x_{r})=\hat{\beta}_{0,r}^{[t]}+\hat{\beta}_{r}^{[t]}x_{r},\qquad\hat{h}_{\beta_{r}}^{[t]}:\mathbb{R}\to\mathbb{R}.$
Evaluating this function at all training observations gives the vector of fitted values
$\hat{\bm{h}}_{\beta_{r}}^{[t]}=\hat{\beta}_{0,r}^{[t]}\bm{1}+\hat{\beta}_{r}^{[t]}\bm{X}_{r},\qquad\hat{\bm{h}}_{\beta_{r}}^{[t]}\in\mathbb{R}^{n},$
where $\bm{1}$ is the $n$-dimensional vector of ones and $\bm{X}_{r}$ is the $n$-dimensional vector containing the $r$-th covariate across all observations.
We define the residual sum of squares for base-learner $r$ in iteration $t$ as
$L_{r}^{[t]}:=\sum_{i=1}^{n}\big(u_{i}^{[t]}-\hat{h}_{\beta_{r}}^{[t]}(x_{i,r})\big)^{2},$
and select the base-learner with minimal loss,
$r^{*}=\arg\min_{r}L_{r}^{[t]}.$
The additive predictor is then updated by moving a small step $\nu$ (typically $\nu=0.1$) in the direction of the selected base-learner. In function notation, this reads
$\hat{\eta}^{[t]}(\bm{x})=\hat{\eta}^{[t-1]}(\bm{x})+\nu\,\hat{h}_{\beta_{r^{*}}}^{[t]}(x_{r^{*}}),\qquad\hat{\eta}^{[t]}:\mathbb{R}^{p}\to\mathbb{R}.$
Equivalently, when evaluated at all training points, the update can be written in vector form as
$\hat{\bm{\eta}}^{[t]}=\hat{\bm{\eta}}^{[t-1]}+\nu\hat{\bm{h}}_{\beta_{r^{*}}}^{[t]},\qquad\hat{\bm{\eta}}^{[t]}\in\mathbb{R}^{n}.$
Note that the same base-learner can be selected multiple times. Consequently, with only one base-learner updated per iteration, a small number of iterations before stopping $t_{stop}$ typically leaves some covariates excluded and, due to the small step length, results in coefficient shrinkage. This makes component-wise gradient boosting conceptually related to penalized regression approaches such as the LASSO and elastic nets [5, 7, 12].
Moreover, unlike black-box machine-learning methods, the resulting additive model remains interpretable and can flexibly accommodate different outcome types—including continuous, binary, count, and time-to-event data—by specifying an appropriate loss function $\rho$. These properties make component-wise boosting particularly attractive for high-dimensional problems that require balancing predictive performance with model sparsity.
The optimal number of iterations, denoted by $t_{stop}^{*}$, is determined by identifying the number of iterations that results in the model with the lowest prediction error. Typically, this number is selected using cross-validation (CV).
2.2 Existing Approaches for Variable Selection with Multiply Imputed Data
Several strategies exist for fitting prediction models on multiply imputed datasets. The simplest is naïve data stacking [13], in which all $M$ imputed datasets are combined into one large dataset comprising $M\times n$ observations. While easy to implement, this approach artificially inflates the sample size, possibly resulting in overfitting.
An alternative approach is estimate averaging [2], wherein model selection is conducted independently within each imputed dataset, and the $M$ resulting sets of coefficient estimates are subsequently averaged. However, as variables selected in one dataset may be excluded in another, the resulting model may contain a substantial number of non‑zero coefficients, thereby compromising model parsimony.
To address this, selection frequency thresholding [3] has been proposed. In this approach, only variables selected in at least a prespecified proportion of imputations are retained, with a 50% cutoff (“majority vote”) being most common. Other thresholds (e.g., 25% or 40%) can be used to balance sparsity and stability. While this approach leads to sparser models than estimate averaging, the cutoff is arbitrary and relevant variables with unstable signals may be excluded. Moreover, dependencies between covariates are partially ignored, and imputation uncertainty is not fully captured.
Beyond these, more advanced approaches have been proposed. Bayesian data‑augmentation methods [14] integrate over the missing‑data distribution within a unified posterior framework, naturally propagating uncertainty, but at the cost of high computational demands and sensitivity to choices of prior distributions. Bootstrapped inclusion‑probability methods [15] use resampling across imputations to stabilize variable selection but require extensive computation and lack clear thresholds for inclusion. EM‑based penalized regression [16] integrates imputation with regularized model fitting, offering efficiency but relying on strong distributional assumptions and custom model‑selection criteria. Finally, joint modeling approaches [17] fit a single model across all imputations by combining likelihoods, although these methods can be complex to implement and are less explored in practice.
Since they deemed existing methods unsatisfactory, Du et al. [4] proposed a novel post‑imputation variable‑selection framework extending regularization methods in two formulations. In the stacked approach, applicable with both LASSO [5] and elastic nets [18], imputations are combined into a single augmented dataset, with each subject’s observations weighted so that their total weight sums to one, yielding a single final model without inflating the sample size. In the grouped approach, applicable only with LASSO, imputed datasets remain independent, resulting in different coefficient estimates, yet coefficients for the same covariate are penalized jointly, enforcing uniform selection across imputations.
In both formulations, the standard and adaptive versions of LASSO [19] and elastic nets [20] can be applied, resulting in a total of six models. However, since Du et al. [4] reported that the stacked, adaptive formulations of LASSO and elastic nets (denoted SaLASSO and SaENET) achieved the best performance, we focus on these two in the comparison with our approach MIBoost. SaLASSO and SaENET, implemented in the R package miselect [6], rely on minimizing a single loss function, like MIBoost.
3 Novel Methods
3.1 The Algorithm MIBoost
The core idea of MIBoost is conceptually simple: although separate gradient boosting algorithms are run on each of the $M$ imputed datasets, base-learner selection at each iteration is based on the aggregated loss, i.e., the sum of the losses across all datasets. This enforces a uniform selection mechanism for base-learners, while gradient computation and coefficient updates remain dataset-specific. The final model is obtained by averaging the fitted base-learner functions across imputed datasets, yielding a single combined predictor.
Algorithm 3.2 provides a mathematical formulation of our procedure. We assume that $M$ imputed datasets are available, each consisting of a column-wise centered design matrix $\bm{X}$ and an outcome vector $y$. Superscripts $(m)$ and $[t]$ denote the imputation index ($m=1,\dots,M$) and the boosting iteration ($t=1,\dots,t_{\mathrm{stop}}$), respectively, where $(m)$ indicates that the corresponding quantity is computed on the $m$-th imputed dataset. Observations with a missing outcome $y$ are assumed to have been excluded prior to imputation, so that $y$ does not depend on $m$. Each imputed dataset $m$ maintains its own additive predictor $\hat{\eta}^{(m)[t]}$, initialized with either the mean of $y^{(m)}$ or zero. The procedure consists of four steps:
  1. (1)
    At the start of each boosting iteration, we compute the negative gradient vector (pseudo-residuals) $\bm{u}^{(m)}$ for each imputed dataset $m$. This vector is obtained as the negative derivative of the loss function with respect to the additive predictor, evaluated at the previous iteration’s predictor $\hat{\bm{\eta}}^{(m)[t-1]}$, indicating the direction of the steepest decrease in the loss.
  2. (2)
    These pseudo-residuals $\bm{u}^{(m)}$ are then regressed on each base-learner $h_{\beta_{r}}$ ($r\in\{1,\dots,p\}$) independently within each imputed dataset, yielding dataset-specific fitted base-learner functions and corresponding residual sums of squares.
  3. (3)
    For each base-learner, the residual sums of squares are summed across the imputed datasets to obtain a single loss value $L_{r}^{[t]}$. The base-learner $r^{*}$ with the lowest aggregated loss in this iteration is then selected. This joint selection step ensures that the same model component is updated in all imputed datasets, improving stability and favoring predictors that are consistently informative across datasets.
  4. (4)
    Finally, the additive predictors $\hat{\bm{\eta}}^{(m)[t]}$ in each imputed dataset are updated by adding the the dataset-specific fitted contribution of the selected base-learner $r^{*}$, scaled by the step length $\nu$ (typically $\nu=0.1$), to the additive predictor from the previous iteration.
(5) Post-processing: The above procedure is repeated for $t_{\mathrm{stop}}$ iterations. After boosting has concluded, the fitted base-learner functions are averaged across imputations to produce a single coherent final additive predictor. Although it is also possible not to pool the models and instead retain $M$ different models, the primary interest usually lies in obtaining a single model. Accordingly, we consider only the pooled model. Algorithm 3.2 summarizes the complete procedure. MIBoost is implemented in the function impu_boost of the R package booami [8], which also provides implementations of estimate averaging and selection-frequency thresholding - as defined in Subsection 2.2 - for direct comparison.
3.2 MIBoostCV - Cross-validation for MIBoost
To select the optimal stopping iteration $t_{\mathrm{stop}}^{*}$ and evaluate out-of-sample performance, $K$-fold cross-validation (CV) can be employed. A detailed mathematical formulation of this procedure is provided in Algorithm 3.2. The steps of the CV procedure are as follows:
Outer loop over CV folds:
  1. (1)
    Split data in training and validation subset: Omit all observations with missing values in the outcome variable $\bm{y}$. The remaining dataset with missing values in the design matrix $\bm{X}$ is first partitioned into $K$ disjoint folds. In each CV iteration, one fold serves as the validation subset, and the remaining K1K-1 folds form the training subset.
  2. (2)
    Multiple imputation: In each CV iteration $k$, the training subset is multiply imputed $M$ times using the chosen imputation procedure, producing $M$ complete training subsets. Crucially, to avoid data leakage, the imputation is performed after the training–validation split. The associated validation subset is then imputed $M$ times using the $M$ imputation models that had been used for the training imputations, ensuring consistency between training and validation imputations. For every imputed training subset, there is thus a corresponding imputed validation subset.
  3. (3)
    Centering: Subsequently, since boosting performs best with standardized inputs, in each of the $M$ imputed training subsets all covariates are centered. The corresponding means from the training subsets are then used to center the covariates in the associated imputed validation subsets.
  4. (4)
    Inner loop over boosting iterations:
    1. (4.1)
      Application of MIBoost: Within each fold, the MIBoost algorithm is applied to the $M$ imputed and centered training subsets for $t_{\mathrm{stop}}$ iterations, with base-learner selection coupled across imputed datasets as described in Algorithm 3.2.
    2. (4.2)
      Average base-learner functions across imputed datasets: At the end of each boosting iteration $t\leq t_{stop}$, the estimated base-learner functions are averaged across the $M$ imputed training subsets, resulting in a single model for that iteration.
    3. (4.3)
      Prediction and error computation: This single model is then used to estimate the response variable $\bm{y}$ in each of the $M$ imputed validation subsets using their respective covariate values. The prediction errors from these $M$ validations are averaged, yielding one aggregated prediction error per boosting iteration $t\leq t_{\mathrm{stop}}$. For each boosting iteration $t$, the averaged prediction error for fold $k$ is stored.
  5. (5)
    Average prediction errors across folds: Since the CV procedure involves $K$ iterations, we obtain $K$ aggregated prediction errors for each boosting iteration. These are then averaged across CV folds to obtain a single final aggregated prediction error for each iteration $t\leq t_{stop}$.
  6. (6)
    Select optimal stopping iteration: Select $t_{\mathrm{stop}}^{*}$ as the boosting iteration that is associated with the smallest final prediction error.
  7. (7)
    Final model fitting: Impute the full dataset with missing values, resulting in $M$ complete training datasets, and center the covariates independently within each imputed dataset. Then, run MIBoost on these datasets for $t_{\mathrm{stop}}^{*}$ boosting iterations to obtain the final model.
The CV procedure is implemented in the R package booami [8]. Specifically, it is available through the function cv_boost_imputed when the user provides already split and imputed data, and through the function cv_boost_raw when starting from a single dataset with missing values. Although described here for MIBoost, this CV framework is generally applicable when performing cross-validation with missing data. In booami, it is implemented not only for MIBoost but also for estimate averaging and selection-frequency thresholding.
Algorithm 1: MIBoost – Multiple-Imputation Gradient Boosting The algorithm requires the following to be specified:
$m\in\{1,\dots,M\}$: index of imputations. $i\in\{1,\dots,n\}$: index of observations. $p$: number of base-learners (candidate covariates). $t_{\mathrm{stop}}$: stopping iteration. $\nu$: step length (typically $\nu=0.1$). $\rho$: differentiable loss function. $\mathcal{H}_{r}$: function space for base-learner $r$; $\hat{h}_{\beta_{r}}^{(m)[t]}\in\mathcal{H}_{r}$ denotes the fitted function at iteration $t$. $\hat{\eta}^{(m)[t]}$: additive predictor function for imputation $m$ at iteration $t$ (with $\hat{\eta}^{(m)[0]}$ as initialization).
  Iterative procedure (for $t=1$ to $t_{\mathrm{stop}}$): 1. Compute negative gradients: For each imputation $m$ and observation $i$, $u_{i}^{(m)[t]}=-\left.\frac{\partial}{\partial\eta_{i}^{(m)}}\rho\!\big(y_{i},\eta_{i}^{(m)}\big)\right|_{\eta_{i}^{(m)}=\hat{\eta}^{(m)[t-1]}(\bm{x}_{i}^{(m)})}.$ 2. Fit base-learners: For each $r\in\{1,\dots,p\}$ and $m$, $\hat{h}_{\beta_{r}}^{(m)[t]}=\arg\min_{h\in\mathcal{H}_{r}}\sum_{i=1}^{n}\big(u_{i}^{(m)[t]}-h(x_{i,r}^{(m)})\big)^{2}.$ 3. Select best base-learner: Compute pooled residual sums of squares and select $r^{*}$ $L_{r}^{[t]}:=\sum_{m=1}^{M}\sum_{i=1}^{n}\Big(u_{i}^{(m)[t]}-\hat{h}_{\beta_{r}}^{(m)[t]}(x_{i,r}^{(m)})\Big)^{2},\qquad r^{*}=\arg\min_{r}L_{r}^{[t]}.$ 4. Update predictors: For each $m$, update the additive predictor function $\hat{\eta}^{(m)[t]}(\bm{x}):=\hat{\eta}^{(m)[t-1]}(\bm{x})+\nu\,\hat{h}_{\beta_{r^{*}}}^{(m)[t]}(x_{r^{*}}).$   5. Post-processing: Average fitted base-learner functions across imputations: $\overline{h}_{\beta_{r}}(x_{r}):=\frac{1}{M}\sum_{m=1}^{M}\hat{h}_{\beta_{r}}^{(m)[t_{\mathrm{stop}}]}(x_{r}),\qquad r=1,\dots,p.$ The final additive predictor is $\hat{\eta}(\bm{x}):=\sum_{r=1}^{p}\overline{h}_{\beta_{r}}(x_{r}).$
Algorithm 2: MIBoostCV The algorithm requires the following to be specified:
$k\in\{1,\dots,K\}$: index of folds. $m\in\{1,\dots,M\}$: index of imputations. $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1,\ldots,n}$: full dataset with $\bm{x}_{i}\in\mathbb{R}^{p\times 1}$. $\mathcal{I}_{k}$: indices of observations in fold $k$. $\mathrm{MIBoost}(\cdot)$: boosting procedure. $t_{\mathrm{stop}}$: maximum number of boosting iterations. $\mathrm{Imp}(\cdot)$: chosen multiple-imputation procedure.
  Outer loop: For each fold $k=1,\dots,K$: 1. Split data in training and validation subset: $\mathcal{D}^{(-k)}=\{(\bm{x}_{i},y_{i}):i\notin\mathcal{I}_{k}\},\quad\mathcal{D}^{(k)}=\{(\bm{x}_{i},y_{i}):i\in\mathcal{I}_{k}\}.$ 2. Multiple imputation: Impute the training subset $M$ times using the chosen imputation model: $\tilde{\mathcal{D}}^{(-k,m)}=\mathrm{Imp}^{(-k,m)}\!\big(\mathcal{D}^{(-k)}\big),\qquad m=1,\dots,M,$ where $\tilde{\mathcal{D}}^{(-k,m)}=\{(\tilde{\bm{x}}_{i}^{(-k,m)},\tilde{y}_{i}):i\notin\mathcal{I}_{k}\}.$ (Tilde $\tilde{\cdot}$ indicates objects containing imputed values.) Apply the same trained $M$ imputation models to the validation subset: $\tilde{\mathcal{D}}^{(k,m)}=\mathrm{Imp}^{(-k,m)}\!\big(\mathcal{D}^{(k)}\big),\qquad m=1,\dots,M,$ with $\tilde{\mathcal{D}}^{(k,m)}=\{(\tilde{\bm{x}}_{i}^{(k,m)},y_{i}):i\in\mathcal{I}_{k}\}.$ 3. Centering (defined on imputed covariates): Compute training-set means for each covariate $j=1,\dots,p$ within each imputed training subset: $\mu_{j}^{(-k,m)}=\frac{1}{|\mathcal{D}^{(-k)}|}\sum_{i\notin\mathcal{I}_{k}}\tilde{x}_{ij}^{(-k,m)}.$ Center the candidate covariates of the training and validation sets component-wise using the training means: $\overline{x}_{ij}^{(-k,m)}=\tilde{x}_{ij}^{(-k,m)}-\mu_{j}^{(-k,m)},\qquad\overline{x}_{ij}^{(k,m)}=\tilde{x}_{ij}^{(k,m)}-\mu_{j}^{(-k,m)},$ for all $i$ and all $j=1,\dots,p$. Define the centered–imputed datasets $\overline{\mathcal{D}}^{(-k,m)}:=\{(\overline{\bm{x}}_{i}^{(-k,m)},y_{i}):i\notin\mathcal{I}_{k}\},\qquad\overline{\mathcal{D}}^{(k,m)}:=\{(\overline{\bm{x}}_{i}^{(k,m)},y_{i}):i\in\mathcal{I}_{k}\}.$ (The bar $\overline{\cdot}$ denotes centered covariates or datasets composed of them. For clarity, the tilde indicating imputation is omitted in these barred symbols; they represent centered versions of the previously imputed values.) 4. Inner loop: Run MIBoost and validate (for $t=1,\dots,t_{\mathrm{stop}}$): 4.1. Apply MIBoost on the imputed and centered training subset: $\{\hat{h}_{\beta_{r}}^{(-k,m)[t]}:r=1,\dots,p\}=\mathrm{MIBoost}\!\left(\overline{\mathcal{D}}^{(-k,m)}\right),$ yielding the fitted base-learner functions at iteration $t$. 4.2. Average base-learners: across fully imputed training subsets (per fold) $\overline{h}_{\beta_{r}}^{(-k)[t]}(\cdot):=\frac{1}{M}\sum_{m=1}^{M}\hat{h}_{\beta_{r}}^{(-k,m)[t]}(\cdot),\qquad r=1,\dots,p.$ (Here and below, an overbar on a function—e.g., $\overline{h}$ or $\overline{\eta}$—denotes averaging across the $M$ imputed training subsets within fold $k$.) Construct the aggregated additive predictor for fold $k$: $\overline{\eta}^{(-k)[t]}(\overline{\bm{x}})=\sum_{r=1}^{p}\overline{h}_{\beta_{r}}^{(-k)[t]}(\overline{x}_{r}),$ with $\overline{h}_{\beta_{r}}^{(-k)[t]}:\mathbb{R}\to\mathbb{R}$ and $\overline{\eta}^{(-k)[t]}:\mathbb{R}^{p}\to\mathbb{R}$, where $\overline{x}_{r}$ is the $r$-th centered covariate in $\overline{\bm{x}}$. 4.3. Prediction and error computation on validation: For each imputed validation subset, predict with $\overline{\eta}^{(-k)[t]}$: $\hat{y}_{i}^{(k,m)}[t]=\overline{\eta}^{(-k)[t]}\!\big(\overline{\bm{x}}_{i}^{(k,m)}\big),\quad i\in\mathcal{I}_{k},$ compute the validation error (per chosen loss) for each $m$, and average these errors across $m$ to obtain the fold-$k$ error at iteration $t$.
  Post-processing across folds: 5. Average prediction errors across folds: For each iteration $t$, average the validation errors across all $K$ folds to obtain $\mathrm{CV\;error}(t)$. 6. Select optimal stopping iteration: Choose $t_{\mathrm{stop}}^{*}=\arg\min_{t\leq t_{\mathrm{stop}}}\mathrm{CV\;error}(t).$ 7. Final model fitting (function notation): Impute the full dataset $M$ times to obtain $\tilde{\mathcal{D}}^{(m)}=\mathrm{Imp}^{(m)}(\mathcal{D})$, center covariates within each imputed dataset using that dataset’s means, and fit MIBoost for $t_{\mathrm{stop}}^{*}$ iterations on each imputed dataset to obtain base-learner functions $\{\hat{h}_{\beta_{r}}^{(m)[t_{\mathrm{stop}}^{*}]}(\cdot)\}_{r=1}^{p}$. Average base-learner functions across imputed datasets and define the final predictor: $\overline{h}_{\beta_{r}}^{[t_{\mathrm{stop}}^{*}]}(x_{r})=\frac{1}{M}\sum_{m=1}^{M}\hat{h}_{\beta_{r}}^{(m)[t_{\mathrm{stop}}^{*}]}(x_{r}),\qquad r=1,\dots,p.$ $\hat{\eta}^{[t_{\mathrm{stop}}^{*}]}(\bm{x}):=\sum_{r=1}^{p}\overline{h}_{\beta_{r}}^{[t_{\mathrm{stop}}^{*}]}(x_{r}).$
4 Simulation Study
4.1 Setup
We compared the performance of our proposed algorithm with established methods for variable selection in multiply imputed datasets. First, we applied traditional component-wise gradient boosting independently to each imputed dataset and derived a single final model using an estimate-averaging approach. We also applied the SaLASSO and SaENET algorithms implemented in the R package miselect. Exact specifications for the latter two are provided in the Appendix.
We performed a simulation study comprising 100 rounds, each generating $n=500$ observations. In each round, the data with missing values were split into a training set (80%) and a test set (20%). Within each training set, all methods applied an additional round of 5-fold cross-validation, repeatedly partitioning the data into inner training and validation subsets to determine the optimal hyperparameters. All computations were performed in R. The following hyperparameters were tuned:
  • number of boosting iterations $t_{\mathrm{stop}}$,
  • shrinkage parameter $\lambda$ in SaLASSO and SaENET,
  • elastic net parameter $\alpha$ in SaENET.
Data‑generating mechanism
The outcome variable was generated using the linear model
$\mathbf{y}=\beta_{0}\mathbf{1}_{n}+\mathbf{X}\bm{\beta}+\bm{\varepsilon},\qquad\beta_{0}=5,\qquad\bm{\varepsilon}\sim\mathcal{N}\!\big(\mathbf{0},5^{2}\mathbf{I}_{n}\big).$
where the design matrix $\mathbf{X}$ contained $p=55$ covariates, $q=5$ of which were informative:
$\beta_{j}\sim\begin{cases}\mathrm{Unif}(1,2),&j\leq 5\\ 0,&j>5.\end{cases}$
We partitioned $\mathbf{X}=[\mathbf{X}_{\mathrm{info}},\mathbf{X}_{\mathrm{noise}}]$, where
$\mathbf{X}_{\mathrm{info}}\sim\mathcal{N}_{q}(\mathbf{0},\Sigma),\quad\Sigma_{ij}=\begin{cases}1&i=j\\ \rho&i\neq j\end{cases},\quad\rho=0.25,\hskip 25.0pt\mathbf{X}_{\mathrm{noise}}\sim\mathcal{N}(0,1).$
Thus, informative covariates ($j\leq 5$) followed a correlated multivariate normal distribution ($\rho=0.25$), while the uninformative covariates ($j>5$) were independently standard normally distributed..
Missingness mechanism and imputation
We introduced missing values in $\mathbf{X}$ under a missing‑at‑random (MAR) mechanism. In addition to $\mathbf{y}$, the first two covariates, $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, were fully observed, with the latter driving the missingness in all other variables ($\mathbf{X}_{j},\;j>2$):
$\mathbb{P}(R_{ij}=0\mid X_{i1},X_{i2})=\mathrm{logit}^{-1}\big(\gamma_{0}+\gamma_{1}X_{i1}+\gamma_{2}X_{i2}\big),$
where $R_{ij}=0$ indicates a missing value, $\bm{\gamma}=(0.75,-0.5)^{\mathsf{T}}$, and the intercept $\gamma_{0}$ was chosen such that the missingness proportion is approximately 30%. Missing values were imputed using multivariate imputation by chained equations (MICE [21]) with predictive mean matching, generating $M=10$ imputed datasets. For each variable, only covariates showing a Spearman correlation of at least 0.1 with that variable were included in the imputation model.
Evaluation metrics
Performance metrics considered were:
  • Mean squared prediction error on test sets
  • Average tuned hyperparameters
  • Mean number of selected variables
  • True positive proportion
  • True negative proportion
The true positive proportion is the proportion of informative covariates included in the final model, while the true negative proportion the proportion of non-informative covariates excluded, both averaged across all simulation rounds. For both measures, values closer to 1 indicate better performance.
4.2 Results
Table 1: Simulation Results
MSPE $\lambda^{*}/\alpha^{*}$ $t_{\mathrm{stop}}^{*}$ TPP TNP # Selec.
EA-Boosting 34.17 113.2 1.00 0.37 36.7
MIBoost 34.11 110.3 1.00 0.78 16.2
SaLASSO 34.40 5.9E-04 1.00 0.76 16.9
SaENET 34.62 5.3E-04/0.74 1.00 0.72 19.1
  • Legend: MSPE = Mean Squared Prediction Error; TPP = True Positive Proportion; TNP = True Negative Proportion; # Selec. = Number of selected variables; EA = Estimate Averaging.
Prediction accuracy was comparable across all evaluated methods (Table 1), with the boosting algorithms performing marginally better overall. In terms of variable-selection performance, all algorithms correctly identified the informative variables but also tended to include uninformative covariates. The model complexities and true negative proportions of MIBoost were similar to those of SaLASSO and SaENET. In contrast, gradient boosting with estimate averaging consistently tended to produced considerably more complex models and included almost two-thirds of the uninformative covariates.
5 Discussion
In our simulation study, MIBoost showed prediction performance similar to that of SaLASSO and SaENET introduced in Du et al. [4], as well as boosting with estimate averaging, but resulted in less complex models than the latter. It is unsurprising that predictive performance remained similar despite these differences in complexity, as accuracy is often only marginally affected by parsimony once the key predictors are included [2]. However, it should be noted that estimate averaging led to the inclusion of substantially more covariates. In practice, the application of such models would require considerably more information, potentially resulting in a more laborious and costly data collection process. Moreover, when model interpretability is a priority, increased complexity inevitably reduces transparency and complicates interpretation.
Furthermore, it is important to emphasize that we introduced MIBoost specifically in the context of component‑wise gradient boosting, which implicitly leads to variable selection, as the need for a single interpretable model derived from multiply imputed datasets is particularly pronounced in statistical learning settings in which interpretability is essential. This choice also facilitates direct comparison with SaLASSO and SaENET. Nonetheless, the underlying principle— selecting, across all imputed datasets, the one base-learner that minimizes an aggregated loss function — can be readily generalized to other boosting frameworks. Accordingly, the approach can be extended in the future to likelihood-based boosting, in which an aggregated likelihood can be defined, as well as to non-statistical boosting methods.
6 Conclusion
In this paper, we proposed MIBoost, a novel approach for applying component-wise gradient boosting to multiply imputed datasets. The method operates by minimizing an aggregated loss function across imputations, ensuring a uniform variable-selection mechanism such that, in every iteration, the same covariate is selected across all datasets. In a simulation study, MIBoost achieved predictive performance comparable to that of other established methods designed to ensure uniform variable selection after multiple imputation. Importantly, the underlying principle of the algorithm can be readily generalized to broader boosting frameworks in future work. To facilitate its application in practice, both MIBoost and its cross-validation framework are available in the open-source R package booami.
Appendix
First of all, it is important to note that the miselect package provides built‑in cross‑validation (CV) functions to find the optimal values for $\lambda$ and $\alpha$. However, these functions can only be applied to data that have already been imputed. As elucidated in subsection 3.2, to prevent data leakage, data should only be imputed after being split into the respective training and validation subsets. We therefore wrote our own code to split the data, perform imputation, and carry out 5‑fold CV.
To find the optimal $\lambda$ and $\alpha$ values, we first needed to define a $\lambda$ sequence for cross-validation. To obtain such a sequence, we initially ran both SaLASSO and SaENET for 50 preliminary simulation rounds using an automatically generated sequence of 200 values for $\lambda$, spaced logarithmically from $\lambda_{\max}$ (the smallest value for which all coefficients are zero) down to $\lambda_{\min}=10^{-6}\lambda_{\max}$. We then recorded the optimal $\lambda$ values obtained across the CV folds. Based on these results, we constructed a new $\lambda$ sequence of length 200 that extended slightly beyond the observed range of optimal values by setting its lower bound to half the smallest selected value and its upper bound to twice the largest selected value. The sequence was again generated on a logarithmic scale, thereby allocating a higher density of values in the region in which most optimal $\lambda$ values were concentrated.
For SaENET only, cross-validation was also used to select an optimal value for $\alpha$. To that end, we created an equally spaced $\alpha$ sequence from 0 to 1, consisting of 41 values.
Having determined the $\lambda$ and $\alpha$ sequences for cross-validation, we ran our custom CV code, which implemented the functions galasso and saenet, but wrapped them in a framework that first split the data into training and validation sets and then performed imputation. For SaLASSO, in each simulation round we selected the $\lambda$ value associated with the smallest CV error. For SaENET, we performed a grid search over all $41\times 200=8200$ $(\lambda,\alpha)$ pairs and selected the pair with the smallest CV error.
Furthermore, the miselect package allows for weighting observations to account for the fact that stacked datasets consist of $M\times n$ observations. For both SaLASSO and SaENET, we chose observation weights of $1/M$, as recommended in [4]. An alternative weighting approach—scaling each observation by the proportion of missing values across all variables—was also discussed in [4]. However, the authors acknowledged that this method would shift the analysis closer to a complete‑case analysis, which runs counter to the purpose of multiple imputation. We therefore did not adopt this approach.
Finally, since we did not apply standard LASSO or elastic nets but but rather their adaptive versions, we needed to define adaptive weight vectors for each run of SaLASSO or SaENET. To do so, we first fitted a preliminary stacked elastic net with equal weights (referred to as SENET in Du et al. [4]). We then extracted the estimated coefficients of all candidate covariates and used them to compute the corresponding adaptive weight vectors.